OpenStudy (anonymous):

Events A and B are overlapping. P(A) = 3/11 P(B) = 5/11 P(A and B) = 2/11 Find P(A or B).

OpenStudy (anonymous):

\[P(A or B) = P(A) + P(B) -P(A and B)\]

OpenStudy (anonymous):

\[P(A U B)= P(A) + P(B) - P(A Intersection B)\]

OpenStudy (anonymous):

what is this this isnt math

OpenStudy (anonymous):

So, \[P(AorB) = \frac{3}{11} + \frac{5}{11}- \frac{2}{11} \implies \frac{6}{11}\]

OpenStudy (anonymous):

tell your teacher to stop asking you fake math

OpenStudy (anonymous):

lolol @ivan514 ,